OpenStudy (redheadangel):

Given the function f(x) = 0.5|x – 4| – 3, for what values of x is f(x) = 7? Yo this is timed so like... THANKS!! WILL FAN AND MEDAL :*

OpenStudy (redheadangel):

x = –24, x = 16 x = –16, x = 24 x = –1, x = 9 x = 1, x = –9 These are the answer choices :)

OpenStudy (redheadangel):

I already tried plugging in the answer choices into the equation, but none of them turned out to answer 7. I just want to know what I'm doing wrong.

OpenStudy (redheadangel):

@pooja195

OpenStudy (haseeb96):

now what you will do we will put 7 at the place of f(x) in the given equation . it will be look like this 7 = 0.5 Ix -4I - 3 7 +3 = 0.5 Ix-4I 10 = 0.5 Ix-4I 10/0.5 = Ix-4I 20 = -(x-4) -x +4 =20 -x = 20 -4 -x = 16 x= -16 is your answer

OpenStudy (redheadangel):

wow thanks! I probably should have thought of that haha, but thanks for the step by step =D

OpenStudy (redheadangel):

would you mind helping me with another? @Haseeb96

OpenStudy (haseeb96):

haha okay i will do but after 10 mins because first i have to help someone else she hired me first. okay

OpenStudy (redheadangel):

Which function is a translation of the parent absolute value function? f(x) = 2|x| f(x) = |x+6| f(x) = |x| f(x) = -4|x|

OpenStudy (redheadangel):

Sure thing!

OpenStudy (redheadangel):

I'm thinking it's either A or C. Probably C, because it came out the most normal looking when I graphed it. lol

OpenStudy (redheadangel):

Also the origin was at (0,0)

OpenStudy (redheadangel):

ooohhh wait it says translation. so this must mean flipped over an axis, but it doesn't say which one?? I'm probably not making any sense, but just wanted to explain my thought process. haha

OpenStudy (redheadangel):

@sleepyhead314 @HelpOfTheGods @Albany_Goon Sorry for tagging random... really need help it's timed haha

OpenStudy (redheadangel):

Which statement is true about -2/3 |x+4-6? The graph of f(x) has a vertex of (–4, 6). The graph of f(x) is horizontally stretched. The graph of f(x) opens upward. The graph of f(x) has a domain of x (less than or equal to) –6.

OpenStudy (redheadangel):

That one is also confusing... I think I just suck at absolute value functions

OpenStudy (redheadangel):

-2/3 |x+4|-6 ***

OpenStudy (redheadangel):

@Owlcoffee @AliceCullen @brunomcv @prowrestler Again, sorry for randomly tagging, ayudame por favor!!

OpenStudy (redheadangel):

Dudes, I got 6 minutes left on this timed thingy. PLEASE. I'll love you forever

OpenStudy (haseeb96):

I think for function , we get C is correct answer

OpenStudy (redheadangel):

okay, that's what i was thinking too thanks, what about the last question?

OpenStudy (haseeb96):

sorry idk for last one

OpenStudy (redheadangel):

that's okay thanks for all your help!!

OpenStudy (haseeb96):

@redheadangel welcome you also